OpenStudy (anonymous):

The first term of an arithmetic sequence is -3 and the fifteenth term is 53. What is the common difference of the sequence? 14/13 25/7 4

OpenStudy (akashdeepdeb):

so a + 14d = 53 a = -3 So put a's value in the equation And get d --> THE COMMON DIFFERENCE!! :D Understood?

OpenStudy (akashdeepdeb):

a + (n-1)d = The nth term! And here n = 15 [RIGHT?] sO 15th term = a + (15-1)d = a+14d Getting me? :D

OpenStudy (jhannybean):

\(a_{15} = 53\) and \(a_1 = -3\)arithmetic sequence : \(\large a_n = a_1 +(n-1)d\) so \[\large 53 = -3 + (15-1)d\]\[\large 53 = -3 + 14d\]\[\large 56 = 14d\]\[\large \frac{56}{14} = d\]